# suppSupplement

### v1

@@ A.3 Scattering matrix
$\hat{\boldsymbol{\xi}}^{\mathrm{out}}=$ $\underbrace{\left[\mathds{1}+\sqrt{\kappa}\left(-\mathrm{i}\sigma_{z}H_{\rm BdG}-\frac{\kappa+\Gamma}{2}\right)^{-1}\sqrt{\kappa}\right]}_{\text{scattering matrix }S}\hat{\boldsymbol{\xi}}^{\mathrm{in}}$ (10a)
$+\underbrace{\left[\sqrt{\kappa}\left(-\mathrm{i}\sigma_{z}H_{\rm BdG}-\frac{\kappa+\Gamma}{2}\right)^{-1}\sqrt{\Gamma}\right]}_{\text{noise matrix}N}\hat{\boldsymbol{\xi}}^{\mathrm{noise}}.$ (10b)
Here, $S$ is the scattering matrix and the matrix $N$ describes the linear response to the fluctuations entering the circuit from the intrinsic loss channels.
In view of discovering classes of solutions displaying the same scattering behavior, we found that it is possible to eliminate the explicit dependence of $S$ and $N$ on the decay rates $\kappa_{i}$ by appropriately rescaling the Bogoliubov de Gennes Hamiltonian and the intrinsic losses. In particular, we find
@@
$N$ $=(S-\mathds{1}_{2N})\sqrt{\gamma}.$ (11b)
with
@@
We note that the rescaled Bogoliubov Hamiltonian $H$ is parametrized by $N$ rescaled detunings $H_{ii}=-\Delta_{i}/\kappa_{i}$ with $1\leq i\leq N$, $N(N-1)/2$ dimensionless couplings $g_{ij}/\sqrt{\kappa_{j}\kappa_{i}}$ with $1\leq i<j\leq N$, and $N(N+1)/2$ squeezing amplitudes $\nu_{ij}/\sqrt{\kappa_{j}\kappa_{i}}$ with $1\leq i\leq j\leq N$. These parameters are learning parameters in the continuous optimization. As pointed out in the main text, the functionality of the device to depends only on dimensionless detunings $H_{ii}$, cooperativities ${\cal C}_{ij}=4|H_{ij}|^{2}$, and the gauge-invariant geometrical phases accumulated on closed loops (synthetic field fluxes), $\Phi_{i,j,l,\ldots,k}=\arg(H_{ij}H_{jl}\ldots H_{ki})$, see Eq. 11a. If we allow for intrinsic loss channels in the port modes, the behaviour is furthermore influenced by the dimensionless intrinsic loss rates $\gamma$, see Eqs. 11a and 11b.
We note that in the most general case, the $S$ matrix is a $2N\times 2N$ matrix, cf Eqs. 4, 11a and 12. This dimensionality reflects the fact that the intensity of the transmitted field can depend on the phase of the input field. For example if both $S_{j,i}\neq 0$ and $S_{j,i+N}\neq 0$ and the input field enters the circuit from port $i$ with average amplitude $\langle a^{\rm in}_{i}\rangle=|\langle a^{\rm in}_{i}\rangle|\exp(i\varphi)$, the intensity $|\langle\hat{a}^{\rm out}_{j}\rangle|^{2}$ of the field transmitted to port $j$ depends on the phase $\varphi$ of the input field, $|\langle\hat{a}^{\rm out}_{j}\rangle|^{2}=|S_{j,i}\exp(i\varphi)+S_{j,i+N}\exp(-i\varphi)|^{2}|\langle a^{\rm in}_{i}\rangle|^{2}$.
To simplify the discussion in the main text we have focused on phase-preserving devices. For such devices, the scattering matrix as defined in Eq. 11a has a block-diagonal form with two $N\times N$ blocks mapped one into the other by the particle-hole symmetry [38]. This allows one to define the scattering matrix $S$ as one of the two blocks obtained from Eq. 11a and, thus, to arrive at an $N\times N$ matrix, consistent with Eq. 1 of the main text. In this cases, our program does not calculate the“large” $2N\times 2N$ matrix but rather only one block.
@@
For phase-preserving amplifiers, we enforce the block structure of the “large” scattering matrix by constraining the matrix $H$ in the following way: We divide the modes into two subsets $M_{1}$ and $M_{2}$ with $N_{1}$ and $N_{2}=N-N_{1}$ modes, respectively. Then, we couple modes in the same set exclusively with beam-splitting interactions, and modes in different sets via two-mode squeezing interactions. The resulting scattering matrix has a block structure because the output fields $a^{\rm out}_{i\in M_{1}}$ and $a^{{\rm out}\dagger}_{i\in M_{2}}$ depend only on the input fields $a^{\rm in}_{i\in M_{1}}$ and $a^{{\rm in}\dagger}_{i\in M_{2}}$, and, likewise, for the fields $a^{\rm out}_{i\in M_{2}}$, $a^{{\rm out}\dagger}_{i\in M_{1}}$, $a^{\rm in}_{i\in M_{2}}$, and $a^{{\rm in}\dagger}_{i\in M_{1}}$. This allows us to take as scattering matrix the block enconding the response of the output fields $a^{\rm out}_{i\in M_{1}}$ and $a^{{\rm out}\dagger}_{i\in M_{2}}$ to the the input fields $a^{\rm in}_{i\in M_{1}}$ and $a^{{\rm in}\dagger}_{i\in M_{2}}$. In practice, we calculate the “small” $N\times N$ scattering matrix by replacing all the $2N\times 2N$ matrices ($\mathds{1}_{2N}$, $\sigma_{z}$, $H$, and $\gamma$) in Eq. 11a with the relevant $N\times N$ block, e.g. we replace $\sigma_{z}$ with
@@ A.4 Stability
In the presence of squeezing interactions, the Langevin equations (5) can describe unstable motion drifting away from a saddle point. The motion is stable as long as the dynamical matrix $D=-\mathrm{i}\sigma_{z}H_{\rm BdG}-(\kappa+\Gamma)/2$ has only eigenvalues with a negative real part.

### v2

@@ Appendix F Scattering matrix
$\hat{\boldsymbol{\xi}}^{\mathrm{out}}=S\hat{\boldsymbol{\xi}}^{\mathrm{in}}+\mathcal{N}\hat{\boldsymbol{\xi}}^{\mathrm{noise}}$ (12)
with
@@
$\mathcal{N}$ $=(S-\mathds{1}_{2N})\sqrt{\gamma}.$ (13b)
Here, $S$ is the scattering matrix and the matrix $\mathcal{N}$ describes the linear response to the fluctuations entering the circuit from the intrinsic loss channels. In view of discovering classes of solutions displaying the same scattering behavior, we found that it is possible to eliminate the explicit dependence of $S$ and $\mathcal{N}$ on the decay rates $\kappa_{i}$ by appropriately rescaling the Bogoliubov de Gennes Hamiltonian and the intrinsic losses by introducing the rescaled Bogoliubov Hamiltonian $H$ as:
@@
We note that $H$ is parametrized by $N$ rescaled detunings $H_{ii}=-\Delta_{i}/\kappa_{i}$ with $1\leq i\leq N$, $N(N-1)/2$ dimensionless couplings $g_{ij}/\sqrt{\kappa_{j}\kappa_{i}}$ with $1\leq i<j\leq N$, and $N(N+1)/2$ squeezing amplitudes $\nu_{ij}/\sqrt{\kappa_{j}\kappa_{i}}$ with $1\leq i\leq j\leq N$. These parameters are learning parameters in the continuous optimization. As pointed out in the main text, the functionality of the device depends only on the dimensionless detunings $H_{ii}$, the cooperativities $C_{ij}=4|H_{ij}|^{2}$, and the gauge-invariant geometrical phases $\Phi_{i,j,l,\ldots,k}=\arg(H_{ij}H_{jl}\ldots H_{ki})$ accumulated in closed loops (synthetic field fluxes), see Eq. 13a. If we allow for intrinsic loss channels in the port modes, the behaviour is furthermore influenced by the dimensionless intrinsic loss rates $\gamma$, see Eqs. 13a and 13b.
We note that in the most general case, the $S$ matrix is a $2N\times 2N$ matrix, cf Eqs. 6, 13a and 14. This dimensionality reflects the fact that the intensity of the transmitted field can depend on the phase of the input field.
To simplify the discussion in the main text we have focused on phase-preserving devices. For such devices, the scattering matrix as defined in Eq. 13a has a block-diagonal form with two $N\times N$ blocks mapped one into the other by the particle-hole symmetry [39]. This allows one to define the scattering matrix $S$ as one of the two blocks obtained from Eq. 13a and, thus, to arrive at an $N\times N$ matrix, consistent with Eq. 3. In this cases, our program does not calculate the“large” $2N\times 2N$ matrix but rather only one block.
@@
For phase-preserving amplifiers, we enforce the block structure of the “large” scattering matrix by constraining the matrix $H$ in the following way: We divide the modes into two subsets $M_{1}$ and $M_{2}$ with $N_{1}$ and $N_{2}=N-N_{1}$ modes, respectively. Then, we couple modes in the same set exclusively with beamsplitter interactions, and modes in different sets via two-mode squeezing interactions. Consequently, the set of equations of motion for $\hat{\boldsymbol{\xi}}=\{\hat{a}_{1},\ldots,\hat{a}_{N},\hat{a}_{1}^{\dagger},\ldots,\hat{a}_{N}^{\dagger}\}$ (see Eq. 7) breaks down into two decoupled sets of equation for $\hat{\tilde{\xi}}$ and $\hat{\tilde{\xi}}^{\dagger}$ which are related with each other via particle-hole symmetry. $\hat{\tilde{\xi}}$ contains all $\hat{a}_{j}$ with $j\in M_{1}$ and all $\hat{a}_{j}^{\dagger}$ with $j\in M_{2}$ and decouples from $\hat{\tilde{\xi}}^{\dagger}$ which contains the respective conjugate operators. Both decoupled sets correspond to the same graph. Our optimisation automatically tests all possible assignments of the modes into the two subsets $M_{1}$ and $M_{2}$.
Given a graph, one can read out which modes are in which subset in the following way: Modes coupled via a beamsplitter interaction are within the same subset, modes coupled via a squeezing interaction are in different subsets. By applying this rule step by step for each coupling, one has characterised all modes quickly.
The resulting scattering matrix has a block structure because the output fields $a^{\rm out}_{i\in M_{1}}$ and $a^{{\rm out}\dagger}_{i\in M_{2}}$ depend only on the input fields $a^{\rm in}_{i\in M_{1}}$ and $a^{{\rm in}\dagger}_{i\in M_{2}}$, and, likewise, for the fields $a^{\rm out}_{i\in M_{2}}$, $a^{{\rm out}\dagger}_{i\in M_{1}}$, $a^{\rm in}_{i\in M_{2}}$, and $a^{{\rm in}\dagger}_{i\in M_{1}}$. This allows us to take as scattering matrix the block enconding the response of the output fields $a^{\rm out}_{i\in M_{1}}$ and $a^{{\rm out}\dagger}_{i\in M_{2}}$ to the the input fields $a^{\rm in}_{i\in M_{1}}$ and $a^{{\rm in}\dagger}_{i\in M_{2}}$. In practice, we calculate the “small” $N\times N$ scattering matrix by replacing all the $2N\times 2N$ matrices ($\mathds{1}_{2N}$, $\sigma_{z}$, $H$, and $\gamma$) in Eq. 13a with the relevant $N\times N$ block, e.g. we replace $\sigma_{z}$ with
@@ Appendix G Stability
In the presence of squeezing interactions, the Langevin equations in Eq. 7 can describe unstable motions drifting away from a saddle point. We can rewrite Eq. 8 in the form:
$\dot{\hat{\boldsymbol{\xi}}}(t)=D\hat{\boldsymbol{\xi}}(t)-\sqrt{\kappa}\hat{\boldsymbol{\xi}}^{\mathrm{in}}(t)-\sqrt{\Gamma}\hat{\boldsymbol{\xi}}^{\mathrm{noise}}(t)$ (16)
with $D=-i\sigma_{z}H_{\rm BdG}-(\kappa+\Gamma)/2$ as the dynamical matrix of the coupled system. The motion is stable as long as the dynamical matrix $D$ has only eigenvalues with a negative real part.